### v1

# Predicting Networks Before They Happen: Experimentation on a Real-Time V2X Digital Twin

Abstract
Emerging safety-critical Vehicle-to-Everything (V2X) applications require networks to proactively adapt to rapid environmental changes rather than merely reacting to them. While Network Digital Twins (NDTs) offer a pathway to such predictive capabilities, existing solutions typically struggle to reconcile high-fidelity physical modeling with strict real-time constraints. This paper presents a novel, end-to-end real-time V2X Digital Twin framework that integrates live mobility tracking with deterministic channel simulation. By coupling the Tokyo Mobility Digital Twin—which provides live sensing and trajectory forecasting—with VaN3Twin—a full-stack simulator with ray tracing—we enable the prediction of network performance before physical events occur. We validate this approach through an experimental proof-of-concept deployed in Tokyo, Japan, featuring connected vehicles operating on 60 GHz links. Our results demonstrate the system’s ability to predict Received Signal Strength (RSSI) with a maximum average error of 1.01 dB and reliably forecast Line-of-Sight (LoS) transitions within a maximum average end-to-end system latency of 250 ms, depending on the ray tracing level of detail. Furthermore, we quantify the fundamental trade-offs between digital model fidelity, computational latency, and trajectory prediction horizons, proving that high-fidelity and predictive digital twins are feasible in real-world urban environments.
I Introduction
Connected and automated vehicles are pushing wireless networks to operate increasingly close to their physical limits, both in terms of reliability and reaction time. Safety critical Vehicle-to-Everything (V2X) applications such as collision avoidance, cooperative perception, and automated intersection management require the network to reason not only on the current state of the environment, but also on how positions and channels will evolve in the immediate future, from tens of milliseconds to a few seconds ahead [11038674]. This is precisely the promise of Network Digital Twins (NDTs): a software replica of the physical system, continuously synchronized with live data, that can be queried to evaluate “what–if” situations before they affect real users [9839640]. Despite significant progress, most existing NDTs operate in an essentially offline fashion. High–fidelity ray tracing, realistic mobility engines, and multi-RAT protocol stacks are routinely used to obtain accurate performance evaluations, but without strict timing constraints, simulation runs may take seconds, minutes, or even hours to complete. This is acceptable for design-time analysis, but incompatible with a reactive or predictive Digital Twin (DT), where results must be delivered within the tight latency budget [9429703] imposed by V2X applications, as an example. If the twin answers too late, vehicles will have moved, the channel will have changed, and any recommended action may already be obsolete or even unsafe. Two recent platforms partially address this gap. The Tokyo Mobility DT—from here on referred to as Mobility DT—developed by the Institute of Science Tokyo, Japan [10443037] provides a detailed 3D reconstruction of the Ookayama Campus area in Tokyo, Japan supporting real or replayed trajectories and enabling realistic urban-scale mobility scenarios. Conversely, VaN3Twin [pegurri2025van3twinmultitechnologyv2xdigital] extends the ms-van3t/ns-3 framework [ms-van3t-journal-2024] by integrating Sionna RT [sionna] in-the-loop, thereby enabling deterministic ray-traced channel prediction and the implementation of Intelligent Transportation Systems (ITS) protocols over full-stack realizations of multiple Radio Access Technologies (RATs). The resulting framework has been validated against on-field measurements, demonstrating close agreement between predicted and observed channel behavior [pegurri2025van3twinmultitechnologyv2xdigital]. Each platform solves a different part of the problem: the Mobility DT offers realism in geometry and mobility, while VaN3Twin offers realism in the complete communication modeling. Neither, however, provides an end-to-end and real-time pipeline capable of ingesting live mobility, predicting near-future trajectories, computing near-future channels, and request action within a strict deadline.
I-A State of the art
General-purpose network simulators such as ns-3 and OMNeT+ form the foundation of most V2X simulation frameworks. Among them, ms-van3t is one of the most comprehensive: it supports IEEE 802.11p, LTE-V2X, NR-V2X, and the full ETSI ITS-G5 facilities layer, and integrates with SUMO and CARLA for mobility simulation [ms-van3t-journal-2024]. Several recent works interpret such simulators as components of broader vehicular NDTs. Examples include blockage-aware mmWave DTs for multi-hop topologies [roongpraiwan2025digital] and multi-modal DTs combining mobility, sensing, and coarse communication abstractions [CazzellaV2XDT]. These platforms move beyond offline simulation but still rely on simplified PHY abstractions to maintain responsiveness. Traditional wireless propagation models in ns-3 and OMNeT++ are stochastic (e.g., 3GPP TR 36.885/38.901) or based on analytical families such as WINNER, able to capture large-scale trends but not the site-specific multipath structure. This limitation becomes critical in DT settings, where small changes in geometry or mobility strongly influence short-term channel evolution. To overcome this, several works integrate external ray-based channel simulators into network simulators. MATLAB-based approaches [matlab-rt-colosseum] emphasize PHY accuracy and hardware emulation, but lack full V2X stacks and scale poorly. Other works import offline ray-traced maps into ns-3 [gaugel2012accurate], enabling bit-level simulation but preventing synchronized updates. OPAL–OMNeT++ integration [RUZNIETO2023100964] achieves realistic geometry but targets non-V2X technologies. VaN3Twin, however, advances this direction by coupling ms-van3t with Sionna RT into a comprehensive V2X NDT framework, enabling high-fidelity, in-the-loop ray-traced propagation and realistic modeling of multi-RAT coexistence over shared spectrum portions. However, despite its channel realism, VaN3Twin has primarily been used offline or in emulation mode, without enforcing a strict real-time prediction deadline. GPU-accelerated ray tracing software such as Sionna RT now enable repeated evaluations with competitive runtime [zhu2024toward]. This has motivated new in-the-loop architectures, including ns3sionna [ns3-sionna-falko] for Wi-Fi and our prior ns-3 + Sionna RT integration [Pegu2505:Toward] for multi-RAT 6G studies. DT-CoVeSS [Twardokus2505:DT-CoVeSS] similarly uses Sionna RT for high-fidelity analysis of C-V2X security. These works demonstrate that in-the-loop ray tracing integration is feasible, but none targets city-scale mobility, full V2X communication stacks, and strict prediction deadlines simultaneously. Vehicular DTs outside ns-3 focus increasingly on near real-time operation. RAVEN [RAVEN-paper] combines live data and contextual information to predict mmWave conditions and support interference management. Vehicle-to-Cloud ADAS [V2C-ADAS] demonstrates cloud-assisted actuation using DT logic. Large-scale mobility DTs such as TuST [TuST-paper-journal] and indoor crowd models [9861008] show that realistic spatio-temporal digital replicas can be built from real data. However, these systems often rely on simplified communication models or separate mobility and communication loops. What is missing is a DT that combines: (i) realistic, city-scale mobility; (ii) a full-stack, ray-traced V2X simulator; and (iii) an explicit timing budget ensuring that predictions are delivered before the corresponding physical time occurs.
Paper Contributions To address the gap identified above, we integrate VaN3Twin within the Tokyo Mobility DT ecosystem using Sionna RT as a shared ray tracing core, orchestrating data exchange and scheduling to enable short-term future channel prediction at latencies compatible with real-time V2X operation. Predictions that arrive too late are discarded to ensure consistency with the evolving physical system. This integration exposes the key challenges of real-time NDT operation, including the computational cost of ray-traced propagation, the synchronization of mobility and geometry between platforms, and the scheduling required to deliver actionable predictions before the corresponding physical time occurs. Rather than proposing new trajectory or channel prediction algorithms, our goal is to quantify how far a real system can be pushed toward real-time predictive operation when combining a live mobility testbed with a high-fidelity V2X DT.
The main contributions of this work are as follows:
  • We design and implement an end-to-end integration of VaN3Twin and the Tokyo Mobility DT into a latency-aware and predictive NDT;
  • We introduce a timing-constrained workflow for future channel state prediction, where mobility updates, 3D scene synchronization, ray tracing, and network simulation are orchestrated to meet a real-time budget;
  • We demonstrate the integrated system in an urban testbed, quantifying the trade-offs between trajectory prediction error, channel prediction accuracy, and system latency.
Paper Organization The remainder of the paper is structured as follows. Section II presents the integration architecture and the real-time workflow. Section III formalizes how model complexity affects computation time and end-to-end system latency. Section IV reports experimental results from the physical deployment in the Ookayama Campus of the Institute of Science Tokyo, Japan. Section V concludes the paper.
II Integration architecture
Figure 1: Schematic architecture of the framework.
This section presents the architecture of the proposed framework and details the interactions among its main components. As shown in Figure 1, the framework integrates three core components: (i) The Physical Entities, i.e., the vehicles equipped with sensing and communication capabilities; (ii) The Mobility DT, which maintains a live digital replica of the surrounding traffic and predicts short-term future mobility states; (iii) VaN3Twin, the NDT engine responsible for full-stack replication and prediction of V2X communication processes. Physical Entities provide time-stamped measurements to the Mobility DT, which produces short-term mobility predictions for VaN3Twin. VaN3Twin uses these predictions to simulate future communication states and returns per-link KPIs. In this architecture, the Mobility DT serves as the coordinating component of the system. It aggregates real-time mobility and perception data collected by LiDAR-equipped Physical Entities as well as by LiDAR and camera sensors mounted on nearby Road Side Units (RSUs) or, eventually, network-enabled positioning techniques [10644093]. Using this information, it continuously maintains and forecasts the mobility state of all dynamic elements in the scenario. Mobility updates are processed at a fixed, pre-defined rate, and predictions are generated for a single, short-term horizon $h$, selected to remain compatible with the end-to-end latency budget of the system. The prediction horizon is kept fixed during operation to ensure deterministic scheduling and bounded computation time. The Mobility DT then orchestrates network-level prediction requests to VaN3Twin and subsequently disseminates predicted communication states back to the Physical Entities—possibly including control commands to be applied to the network configuration. Communication between system elements is structured according to their functional roles. A specific physical entity provides the interface between Physical Entities and the Mobility DT, leveraging a 2.4 GHz IEEE 802.11n Wi-Fi Control Plane (CP) for low latency over-the-air control operations. Meanwhile, the Mobility DT and VaN3Twin communicate over a wired Ethernet Local Area Network (LAN), reflecting their co-location at the network edge to minimize latency. All inter-component exchanges employ lightweight UDP messages with JSON-formatted payloads, enabling efficient and low-overhead transmission of state and control information. Message delivery is best-effort: occasional packet losses are tolerated and naturally overridden by subsequent updates, avoiding retransmission delays. All components rely on a shared time base, with timestamps carried in each message and clocks synchronized using standard network time synchronization mechanisms, ensuring consistent temporal alignment across mobility and communication predictions. The detailed data exchange loop is described in the following subsection.
II-A End-to-end Workflow
Figure 2: End-to-end communication workflow.
The end-to-end workflow illustrated in Figure 2 specifies all the interactions introduced above and clarifies how mobility prediction and communication prediction are coupled in practice. Physical Entities periodically transmit their sensed location and measured communication data to the Mobility DT—located in the edge cloud—with an implementation-dependent reporting period $\Delta t_{\mathrm{PE}}$. These updates are time-stamped and provide the input required to keep the digital replica synchronized with the physical system and to continuously adjust the internal prediction models. Based on the aggregated and synchronized mobility state, the Mobility DT computes short-term trajectory forecasts for all relevant dynamic objects. Network prediction requests toward VaN3Twin are then generated according to a configurable policy, which may be periodic or event-driven, and always target a single future time instant $t+h$. The prediction horizon $h$ is chosen so as to remain compatible with the end-to-end latency budget; multi-horizon requests are also possible in principle but are outside the scope of this work. Upon receiving a request, VaN3Twin instantiates a future communication scenario corresponding to the target time $t+h$. Predicted vehicle poses are used to construct a consistent 3D scene, which serves as input to the deterministic ray-based channel model formalized in Section III. The resulting channel realizations drive full-stack V2X simulations, possibly yielding protocol- and packet-level performance indicators. In parallel, the recently observed communication measurements are used to compensate for systematic prediction biases, such as offsets on the computed received power. The predicted communication state is then returned to the Mobility DT and made available for network-level decision making. When required, the DT translates these predictions into control actions which are conveyed back to the Physical Entities over the dedicated Wi-Fi CP, thereby closing the loop between the digital and physical domains. To reason about the limits of this approach, the next section provides a formal characterization of the proposed integration, detailing the underlying trajectory prediction model, the channel estimation process, and the resulting trade-offs between model complexity, end-to-end latency, and the maximum achievable prediction horizon.
III End-to-End Analysis
This section provides a formal characterization of the integration presented in Section II. We describe the trajectory prediction model, formalize the ray-based channel prediction process, and introduce a compact notion of model complexity that links physical-model fidelity to end-to-end latency and feasible prediction horizons.
III-A Mobility and Channel Prediction
Each vehicle $i$ in the scenario is represented by the kinematic state $\mathbf{x}^{i}_{t}=(\mathbf{p}^{i}_{t},\,\mathbf{v}^{i}_{t},\,\psi^{i}_{t})$. The state evolution assumes a constant-velocity motion model, with $\mathbf{F}^{i}_{t}$ defined accordingly based on the sampling interval. $\mathbf{z}^{i}_{t}$ represents the state measurement vector, which the DT fuses into a unified state observation. The temporal evolution of each vehicle follows a linear state-space model defined by the state-transition matrix $\mathbf{F}^{i}_{t}$, the control-input matrix $\mathbf{B}^{i}_{t}$, and the measurement model matrix $\mathbf{M}^{i}_{t}$, with process noise covariance $\mathbf{Q}^{i}_{t}$ and measurement noise covariance $\mathbf{R}^{i}_{t}$. The process noise covariance $\mathbf{Q}^{i}_{t}$ accounts for deviations from ideal constant-velocity motion and is treated as a design parameter, whereas the measurement noise covariance $\mathbf{R}^{i}_{t}$ is sensor-dependent and reflects the heterogeneous accuracy of the contributing sensing modalities. Given the filtered estimate at time $t$, the DT propagates the state $h$ steps forward to obtain the predicted pose $\hat{\mathbf{x}}^{\,i}_{t+h|t}$, which defines the forecasted traffic geometry used for channel prediction (Algorithm 1).
Algorithm 1 Trajectory Prediction for Vehicle $i$
[…]
1: Prediction:
[…]
4: Update:
[…]
8: Horizon propagation:
[…]
10:Return: predicted pose $\hat{\mathbf{x}}^{\,i}_{t+h|t}$
Once the future scene configuration is assembled, the DT evaluates the wireless channel through a deterministic ray-based propagation model. For each transmitter–receiver pair $(i,j)$, the ray tracing engine emits $N_{\mathrm{rays}}$ rays from the antenna of vehicle $i$, with departure directions covering the unit sphere. Each ray propagates through the environment and may undergo up to $N_{\mathrm{int}}$ interactions with surrounding surfaces. A ray reaching the antenna of vehicle $j$ constitutes a valid propagation path and is added to the set $\mathcal{P}^{i,j}_{\mathrm{valid}}$. Each path $p\in\mathcal{P}^{i,j}_{\mathrm{valid}}$ is characterized by its length $d_{p}^{i,j}$, delay $\tau_{p}^{i,j}=d_{p}^{i,j}/c$, angles of departure $(\theta^{(p)}_{\mathrm{T}},\varphi^{(p)}_{\mathrm{T}})$ and arrival $(\theta^{(p)}_{\mathrm{R}},\varphi^{(p)}_{\mathrm{R}})$, and by an ordered sequence of $N_{p}$ interactions. Its effect on the received field is described by the $2\times 2$ propagation matrix:
$\mathbf{A}_{p}^{i,j}=\left(\frac{\lambda}{4\pi d_{p}^{i,j}}\right)e^{-j\frac{2\pi}{\lambda}d_{p}^{i,j}}\prod_{\ell=1}^{N_{p}}\mathbf{M}_{p}^{(\ell)},$ (1)
where each $\mathbf{M}_{p}^{(\ell)}$ models the polarization-dependent transformation induced by the $\ell$-th interaction. The transmit and receive antennas are represented through their complex pattern vectors $\mathbf{c}_{\mathrm{T}}(\theta,\varphi)$ and $\mathbf{c}_{\mathrm{R}}(\theta,\varphi)$. The scalar baseband contribution of path $p$ to the channel between vehicles $i$ and $j$ is then
$g_{p}^{i,j}=\mathbf{c}^{\mathsf{H}}_{\mathrm{R}}\!\left(\theta^{(p)}_{\mathrm{R}},\varphi^{(p)}_{\mathrm{R}}\right)\mathbf{A}_{p}^{i,j}\mathbf{c}_{\mathrm{T}}\!\left(\theta^{(p)}_{\mathrm{T}},\varphi^{(p)}_{\mathrm{T}}\right)$. (2)
The predicted multipath channel at time $t+h$ is therefore
$h^{i,j}_{t+h|t}(\tau)=\sum_{p\in\mathcal{P}^{i,j}_{\mathrm{valid}}}g_{p}^{i,j}\,\delta\!\left(\tau-\tau_{p}^{i,j}\right).$ (3)
Collecting all pairwise impulse responses yields the predicted channel matrix $\widehat{\mathbf{H}}_{t+h|t}=\big(h^{i,j}_{t+h|t}\big)_{i,j\in\mathcal{V}}.$
III-B Computational Complexity and Execution Time
TABLE I: Ray tracing parameters for each DI.
RT Parameter DI 1 DI 2 DI 3 DI 4 DI 5
#\# Max. Interactions 3 3 5 8 8
#\# Rays per source $10^{3}$ $10^{3}$ $^{6} 1010$$^{10}$ $10^{10}$
Direct LoS Path $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Specular reflection $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Diffuse reflection $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
Refraction $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
Diffraction $\times$ $\times$ $\times$ $\times$ $\checkmark$
Let $\tau_{e2e}$ be the total time elapsed between the acquisition of a measurement by a Physical Entity and the completion of the corresponding DT update. Formally:
$\tau_{e2e}=\tau_{m}+\tau_{w}+\tau_{tp}+\tau_{req}+\tau_{rt}+\tau_{req}+\tau_{w},$ (4)
where $\tau_{m}$ is the measurement acquisition time, $\tau_{w}$ the communication delay between vehicles and the Mobility DT, $\tau_{tp}$ the Mobility DT update and trajectory prediction time, $\tau_{req}$ the inter-process communication delay, and $\tau_{rt}$ the execution time of the physical-modeling stage. Given the computational complexity of the ray tracing operation, $\tau_{rt}$ is the dominant and most variable value. Specifically, it depends on the level of physical detail included in the propagation model. To describe these configurations in a compact and general manner, we introduce the Ray Tracing Detail Index, denoted DI, a scalar metric used to classify the fidelity level of the physical-modeling stage. Each DI level corresponds to a specific configuration of the physical model, ranging from lightweight approximations to high-fidelity reconstructions. Table I reports the five DI levels considered in this work. Lower levels include only the essential components needed for coarse propagation estimates, whereas higher levels progressively incorporate more detailed interaction mechanisms, finer spatial resolution, and deeper multipath exploration. Although in our implementation these fidelity levels map to different configurations of a ray-based propagation model, the abstraction provided by DI remains general and applicable to any physical-modeling backend supporting multiple complexity tiers. As the DI increases, the execution time of the physical model, denoted $\tau_{rt}$, grows accordingly. To ensure causal consistency, the prediction horizon $h$ must exceed the end-to-end latency $\tau_{e2e}$. Increasing model fidelity therefore requires larger horizons, which in turn amplify trajectory prediction errors and their impact on channel estimation, revealing a fundamental trade-off between physical-model complexity, latency, and predictive accuracy. Section IV provides an empirical characterization of these trade-offs based on real-world experiments conducted in Tokyo, Japan, evaluating how different DI levels impact both computational latency and channel prediction accuracy.
IV Experimental and Simulated Results
To demonstrate the practical operation of the proposed system, we consider a proof-of-concept deployment carried out in the Ookayama campus of the Institute of Science Tokyo, Japan complete with measured end-to-end latencies. To complement the experimental analysis, we then consider a simulation study performed in the identical geometric scenario in which we systematically vary the DI and introduce synthetic perturbations into the trajectory prediction module.
IV-A Scenario introduction
(a) Real picture of the scenario.
(b) VaN3Twin scenario with trajectory.
(c) Example of ray-traced propagation rays.
Figure 3: Real picture (a), the corresponding VaN3Twin ray tracing representation (b), and an example of the computed propagation rays (c) of the road segment used in the experiments on the Ookayama campus of the Institute of Science Tokyo, Japan.
TABLE II: VaN3Twin settings for ray tracing.
RT Settings Value
Carrier Frequency 60 GHz
Buildings and Roads Material Concrete [iturp2040]
Trees Material Wood [iturp2040]
Cars Material Metal [iturp2040]
Type of Antennas Panasonic WiGig RSU
Polarization Vertical
Digital Model DI DI 5 (see Tab. I)
The evaluation scenario involves three connected vehicles (Figure 3(a)) that interact with the Mobility DT through a $2.4$ GHz Wi-Fi CP, and are each equipped with a roof-mounted Panasonic WiGig RSU forming the $60$ GHz User Plane (UP). The Mobility DT and the VaN3Twin engine are deployed at the edge—corresponding to the location of Vehicle 3—and exchange data through a wired Ethernet LAN. Figure 3(b) illustrates the corresponding VaN3Twin ray tracing scenario of the considered street in Ookayama, Tokyo, Japan. Buildings, vehicles, and antennas are modeled in VaN3Twin using material properties and radiation patterns consistent with the experimental setup (see Table II). VaN3Twin ray tracing details are reported in Table II. During the experimental campaign, the vehicles perform iperf-based data transmissions over the WiGig UP. Vehicle 1 (a Toyota COMS) and Vehicle 3 (a motorized wheel cart) remain stationary and maintain a direct Line of Sight (LoS) link, while Vehicle 2 (a Toyota Estima) follows a predefined trajectory (Figure 3(b)) that intermittently obstructs the link. The Mobility DT is used to predict, in real time, the $60$ GHz channel between Vehicles 1 and 3 via VaN3Twin, based on the predicted trajectory of Vehicle 2 with prediction horizon $h=500$ ms. The Mobility DT simultaneously collects from each vehicle its current position and the field-measured Received Signal Strength Indicators (RSSIs), which are employed for in-the-loop calibration of the prediction process. After computation, the predicted RSSI values are transmitted to the vehicles through the Wi-Fi CP. To quantify the effectiveness of this real-time predictive loop, we evaluate the accuracy of the RSSI forecasts produced by VaN3Twin along the trajectory of Vehicle 2. The Mobility DT employs the simple predictor for short-term trajectory estimation presented in Section III-A, while VaN3Twin operates using the DI 5 configuration reported in Table I. Results show a maximum average RSSI prediction error of $1.01$ dB and a 95th-percentile error of $3.05$ dB, demonstrating accurate prediction of the $60$ GHz link dynamics under mobility-induced obstruction. Since VaN3Twin operates on predicted positions, channel accuracy is jointly limited by trajectory error and by the ray tracing DI, which directly determines the end-to-end latency and thus the minimum prediction horizon required for real-time operation.
IV-B End-to-End System Latency
Building on the formal definition introduced in Section III-B, this section reports on-field measurements of the end-to-end latency $\tau_{\mathrm{e2e}}$, decomposed according to (4). All the measured fixed latency components are summarized in Table III, while the resulting end-to-end latency distributions as a function of the DI level are shown in Figure 4.
Fixed system latency components: measurement acquisition ($\tau_{m}$) was evaluated by collecting location and channel samples over $60$ s with a $100$ ms sampling interval, yielding approximately $600$ samples per experiment. No statistically significant differences were observed between static and dynamic conditions. Communication-related delays ($\tau_{w}$ and $\tau_{req}$) were measured using ICMP-based Round Trip Time (RTT) probing over $100$ messages and conservatively converted to one-way latency by halving the resulting value. Mobility DT update time and trajectory prediction latency ($\tau_{tp}$) was obtained by profiling the time from the scenario update up to the execution of the linear prediction algorithm described in Section III-A. These components—resumed in Table III—constitute a fixed system latency budget, independent of the DI configuration.
Ray tracing execution times ($\tau_{rt}$): were measured for the five DI configurations defined in Table I using VaN3Twin with Sionna RT v1.2.1. Experiments were conducted on a workstation with an NVIDIA GeForce RTX 3090 GPU and Intel Core i9-10980XE CPUs running Ubuntu 22.04.5. $\tau_{rt}$ increases monotonically with the DI level, reflecting the higher computational load induced by denser ray interactions and more complex propagation mechanisms.
Figure 4: Distribution of the measured end-to-end latency $\tau_{e2e}$ for different Ray Tracing DI configurations.
TABLE III: Fixed latency components contributing to $\tau_{e2e}$.
Component Resulting average
Measurement acquisition $\tau_{m}$ $8.9$ ms
Wi-Fi CP communication $\tau_{w}$ $1.1$ ms
Trajectory prediction $\tau_{tp}$ $4.4$ ms
DT–VaN3Twin communication $\tau_{req}$ $0.6$ ms
Total fixed latency $\mathbf{16.2}$ ms
End-to-end latency $\tau_{\mathrm{e2e}}$: is obtained by combining the fixed system latency components with the DI-dependent ray tracing execution time. The resulting distributions of $\tau_{\mathrm{e2e}}$ for each DI configuration are reported in Figure 4, highlighting the progressive increase in both latency and variability as the ray tracing complexity grows. These results demonstrate that improving channel modeling fidelity by increasing the DI level inevitably enlarges the end-to-end latency budget. Since channel prediction must be available before transmission, this latency directly translates into a minimum required trajectory prediction horizon. As a result, higher-fidelity channel modeling forces the DT to rely on longer-term mobility forecasts.
IV-C Impact of trajectory prediction error
Because the end-to-end latency of the DT determines how far into the future vehicle positions must be predicted, longer latencies inevitably amplify trajectory prediction errors. This section quantifies how such latency-induced positional inaccuracies propagate into ray-traced channel prediction.
Let $\mathbf{p}(t)=[x(t),\,y(t)]^{\top}$ denote the ground-truth position at time $t$, and let $\tilde{\mathbf{p}}(t)=\mathbf{p}(t)+\mathbf{e}(t)$, be its perturbed counterpart, where $\mathbf{e}(t)$ models the prediction error. A normalized parameter $k\in[0,1]$ controls the maximum error magnitude through $\varepsilon_{k}=k\,\varepsilon_{\max}$, with $\varepsilon_{\max}=1$ m in the considered setup. The displacement vector is sampled uniformly from a circular annulus $\mathbf{e}(t)\sim\mathcal{A}\!\left(\tfrac{1}{3}\varepsilon_{k},\;\tfrac{1}{2}\varepsilon_{k}\right),$ and $\mathbf{e}(t)$ is isotropically sampled within the annulus itself. To quantify the impact of trajectory prediction inaccuracies, we evaluate two complementary metrics: (i) the deviation in received signal strength and (ii) the stability of the geometric visibility state. Both metrics are computed by comparing the channels obtained from the perturbed positions $\tilde{\mathbf{p}}(t)$ with those obtained from the true positions $\mathbf{p}(t)$.
Figure 5: Experimental RSSI RMSE (blue) and LoS Prediction Accuracy (green) trends for different trajectory prediction perturbation.
RSSI Prediction Error: let $\mathrm{RSSI}(t)$ denote the RSSI computed from $\mathbf{p}(t)$, and let $\widetilde{\mathrm{RSSI}}(t)$ denote the value obtained from $\tilde{\mathbf{p}}(t)$. The discrepancy induced by a perturbation index $k$ is measured through the Root Mean Squared Error (RMSE):
$\mathrm{RMSE}_{k}=\sqrt{\frac{1}{N}\sum_{t=1}^{N}\left(\widetilde{\mathrm{RSSI}}(t)-\mathrm{RSSI}(t)\right)^{2}},$ (5)
where $N$ is the number of channel realizations. Figure 5 shows the dependence of the computed RMSE (blue) on the perturbation level $k$, focusing on measurements collected during the LoS to non-LoS transition caused by the obstructing vehicle (Vehicle 2). The relationship is clearly non-linear, with a rapid error increase for small positional deviations. For the considered scenario, $k\approx 0.1$ already yields an RMSE of about $20$ dB on RSSI prediction. Since small geometric inaccuracies already induce large prediction errors, these results also anticipate the sensitivity of the LoS classification, analyzed next.
LoS Classification Robustness: trajectory perturbations may also change the geometric visibility between a transmitter and a receiver. Let $\mathrm{LoS}_{\mathrm{ref}}(t)$ denote the visibility classification obtained from $\mathbf{p}(t)$ and $\widetilde{\mathrm{LoS}}(t)$ that obtained from $\tilde{\mathbf{p}}(t)$. The robustness of the visibility classification under perturbation level $k$ is expressed by
$\eta_{k}=\frac{\mathrm{TP}_{k}+\mathrm{TN}_{k}}{N},$ (6)
where $\mathrm{TP}_{k}$ counts the instants for which both classifications report a LoS condition and $\mathrm{TN}_{k}$ counts the instants for which both report a non-LoS condition. The metric $\eta_{k}$ therefore quantifies the agreement between the reference and perturbed LoS classifications under a positional perturbation of magnitude $k$. Figure 5 shows $\eta_{k}$ (green) as a function of $k$. The geometric visibility prediction accuracy decreases monotonically, indicating a steady degradation of classification reliability with increasing positional error. For $k<0.1$, the accuracy remains close to $100\%$, whereas it degrades substantially as $k$ increases. At the maximum considered error, $\varepsilon_{k}\approx 1$ m ($k=1$), the consistency drops to about $75\%$.
In conclusion, positional errors induced by longer prediction horizons offset the gains of high-fidelity ray tracing, exposing a fundamental trade-off between channel modeling precision and end-to-end latency.
V Conclusions
In this work, we presented and experimentally validated a real-time, predictive NDT for V2X communications, explicitly designed around end-to-end latency constraints. By integrating a live Mobility Digital Twin with the VaN3Twin engine, we implemented a closed-loop workflow capable of operating in synchrony with the physical environment. A key outcome of this study is the end-to-end latency analysis of the Digital Twin pipeline, covering sensing, mobility prediction, ray-tracing-based channel computation, and data exchange. This analysis allowed us to identify how the computational complexity of the physical model directly constrains the feasible prediction horizon and, consequently, the usefulness of the Digital Twin for real-time operation. Experimental results obtained in a live urban environment demonstrate that high-fidelity channel modeling can be executed within a bounded time budget, provided that model complexity and scheduling are jointly designed. Overall, this work shows that real-time integration of mobility Digital Twins and ray-tracing-based network modeling is feasible and practically deployable. Future work will extend this latency-aware framework toward real-time prediction of higher-layer network performance metrics leveraging on VaN3Twin full-stack V2X simulation capability.
Acknowledgment
This work was partially supported by the European Union - Next Generation EU under the Italian National Recovery and Resilience Plan (NRRP), Mission 4, Component 2, Investment 1.3,CUP D43C22003080001, partnership on “Telecommunications of the Future” (PE00000001 - program “RESTART”).

### v2

@@
 Connected and automated vehicles are pushing wireless networks to operate increasingly close to their physical limits, both in terms of reliability and reaction time.
-Safety critical Vehicle-to-Everything (V2X) applications such as collision avoidance, cooperative perception, and automated intersection management require the network to reason not only on the current state of the environment, but also on how positions and channels will evolve in the immediate future, from tens of milliseconds to a few seconds ahead [11038674].
+Safety critical Vehicle-to-Everything (V2X) applications such as collision avoidance, cooperative perception, and automated intersection management require the network to reason not only about the current state of the environment, but also on how positions and channels will evolve in the immediate future, from tens of milliseconds to a few seconds ahead [4].
 This is precisely the promise of Network Digital Twins (NDTs): a software replica of the physical system, continuously synchronized with live data, that can be queried to evaluate “what–if” situations before they affect real users [6].
-Despite significant progress, most existing NDTs operate in an essentially offline fashion.
+Despite significant progress, most existing NDTs operate in an essentially offline manner.
 High–fidelity ray tracing, realistic mobility engines, and multi-RAT protocol stacks are routinely used to obtain accurate performance evaluations, but without strict timing constraints, simulation runs may take seconds, minutes, or even hours to complete.
-This is acceptable for design-time analysis, but incompatible with a reactive or predictive Digital Twin (DT), where results must be delivered within the tight latency budget [9429703] imposed by V2X applications, as an example.
-If the twin answers too late, vehicles will have moved, the channel will have changed, and any recommended action may already be obsolete or even unsafe.
+This is acceptable for design-time analysis, but incompatible with an online, predictive Digital Twin (DT), for which results must be delivered within the tight latency budget [20], for example as imposed by V2X applications.
+If the twin delivers its prediction too late, vehicles will have moved, the channel will have changed, and any recommended action may already be obsolete or even unsafe.
 Two recent platforms partially address this gap.
-The Tokyo Mobility DT—from here on referred to as Mobility DT—developed by the Institute of Science Tokyo, Japan [10443037] provides a detailed 3D reconstruction of the Ookayama Campus area in Tokyo, Japan supporting real or replayed trajectories and enabling realistic urban-scale mobility scenarios.
-Conversely, VaN3Twin [pegurri2025van3twinmultitechnologyv2xdigital] extends the ms-van3t/ns-3 framework [ms-van3t-journal-2024] by integrating Sionna RT [sionna] in-the-loop, thereby enabling deterministic ray-traced channel prediction and the implementation of Intelligent Transportation Systems (ITS) protocols over full-stack realizations of multiple Radio Access Technologies (RATs).
-The resulting framework has been validated against on-field measurements, demonstrating close agreement between predicted and observed channel behavior [pegurri2025van3twinmultitechnologyv2xdigital].
-Each platform solves a different part of the problem: the Mobility DT offers realism in geometry and mobility, while VaN3Twin offers realism in the complete communication modeling.
-Neither, however, provides an end-to-end and real-time pipeline capable of ingesting live mobility, predicting near-future trajectories, computing near-future channels, and request action within a strict deadline.
+The Tokyo Mobility DT—from here on referred to as Mobility DT—developed by the Institute of Science Tokyo, Japan [18] provides a detailed 3D reconstruction of the Ookayama Campus area in Tokyo, Japan.
+It supports both real-time and replayed trajectories enabling realistic urban-scale mobility scenarios.
+Conversely, VaN3Twin [9] extends the ms-van3t/ns-3 framework [13] with the in-the-loop integration of Sionna RT [5], thereby enabling deterministic ray-based channel prediction and the implementation of Intelligent Transportation Systems (ITS) protocols over full-stack realizations of multiple Radio Access Technologies (RATs).
+The resulting framework has been validated against on-field measurements, demonstrating close agreement between predicted and observed channel behavior [9].
+Each platform solves a different part of the problem: the Mobility DT offers realism in geometry and mobility, while VaN3Twin offers realism across the complete communication stack.
+Neither, however, provides an end-to-end real-time pipeline capable of ingesting live mobility, predicting near-future trajectories, computing near-future channels, and take action within a strict deadline.
 I-A State of the art
-General-purpose network simulators such as ns-3 and OMNeT+ form the foundation of most V2X simulation frameworks.
-Among them, ms-van3t is one of the most comprehensive: it supports IEEE 802.11p, LTE-V2X, NR-V2X, and the full ETSI ITS-G5 facilities layer, and integrates with SUMO and CARLA for mobility simulation [ms-van3t-journal-2024].
+General-purpose network simulators such as ns-3 and OMNeT++ form the foundation of most V2X simulation frameworks.
+Among them, ms-van3t is one of the most comprehensive: it supports IEEE 802.11p, LTE-V2X, NR-V2X, and the full ETSI ITS-G5 facilities layer, and integrates with SUMO and CARLA for mobility simulation [13].
 Several recent works interpret such simulators as components of broader vehicular NDTs.
-Examples include blockage-aware mmWave DTs for multi-hop topologies [roongpraiwan2025digital] and multi-modal DTs combining mobility, sensing, and coarse communication abstractions [CazzellaV2XDT].
+Examples include blockage-aware mmWave DTs for multi-hop topologies [14] and multi-modal DTs combining mobility, sensing, and coarse communication abstractions [1].
 These platforms move beyond offline simulation but still rely on simplified PHY abstractions to maintain responsiveness.
@@
 To overcome this, several works integrate external ray-based channel simulators into network simulators.
-MATLAB-based approaches [matlab-rt-colosseum] emphasize PHY accuracy and hardware emulation, but lack full V2X stacks and scale poorly.
-Other works import offline ray-traced maps into ns-3 [gaugel2012accurate], enabling bit-level simulation but preventing synchronized updates.
-OPAL–OMNeT++ integration [RUZNIETO2023100964] achieves realistic geometry but targets non-V2X technologies.
+MATLAB-based approaches [15] emphasize PHY accuracy and hardware emulation, but lack full V2X stacks and exhibit poor scalability.
+Other works import offline ray-traced maps into ns-3 [3], enabling bit-level simulation but preventing synchronous updates.
+OPAL–OMNeT++ integration [16] achieves realistic geometry but targets non-V2X technologies.
 VaN3Twin, however, advances this direction by coupling ms-van3t with Sionna RT into a comprehensive V2X NDT framework, enabling high-fidelity, in-the-loop ray-traced propagation and realistic modeling of multi-RAT coexistence over shared spectrum portions.
 However, despite its channel realism, VaN3Twin has primarily been used offline or in emulation mode, without enforcing a strict real-time prediction deadline.
-GPU-accelerated ray tracing software such as Sionna RT now enable repeated evaluations with competitive runtime [zhu2024toward].
-This has motivated new in-the-loop architectures, including ns3sionna [ns3-sionna-falko] for Wi-Fi and our prior ns-3 + Sionna RT integration [Pegu2505:Toward] for multi-RAT 6G studies.
-DT-CoVeSS [Twardokus2505:DT-CoVeSS] similarly uses Sionna RT for high-fidelity analysis of C-V2X security.
+GPU-accelerated ray tracing software such as Sionna RT now enable repeated evaluations with competitive runtime [21].
+This has motivated new in-the-loop architectures, including ns3sionna [22] for Wi-Fi and our prior ns-3 + Sionna RT integration [10] for multi-RAT 6G studies.
+DT-CoVeSS [17] similarly uses Sionna RT for high-fidelity analysis of CV2X security.
 These works demonstrate that in-the-loop ray tracing integration is feasible, but none targets city-scale mobility, full V2X communication stacks, and strict prediction deadlines simultaneously.
 Vehicular DTs outside ns-3 focus increasingly on near real-time operation.
-RAVEN [RAVEN-paper] combines live data and contextual information to predict mmWave conditions and support interference management.
-Vehicle-to-Cloud ADAS [V2C-ADAS] demonstrates cloud-assisted actuation using DT logic.
-Large-scale mobility DTs such as TuST [TuST-paper-journal] and indoor crowd models [9861008] show that realistic spatio-temporal digital replicas can be built from real data.
+RAVEN [2] combines live data and contextual information to predict mmWave conditions and support interference management.
+Vehicle-to-Cloud ADAS [19] demonstrates cloud-assisted actuation using DT logic.
+Large-scale mobility DTs such as TuST [11] and indoor crowd models [12] show that realistic spatio-temporal digital replicas can be built from real data.
 However, these systems often rely on simplified communication models or separate mobility and communication loops.
-What is missing is a DT that combines: (i) realistic, city-scale mobility; (ii) a full-stack, ray-traced V2X simulator; and (iii) an explicit timing budget ensuring that predictions are delivered before the corresponding physical time occurs.
+What is missing is a DT that combines: (i) realistic, city-scale mobility; (ii) a full-stack, ray-traced V2X simulator; and (iii) an explicit timing budget ensuring that predictions are delivered before the event takes place in the physical world.
 Paper Contributions To address the gap identified above, we integrate VaN3Twin within the Tokyo Mobility DT ecosystem using Sionna RT as a shared ray tracing core, orchestrating data exchange and scheduling to enable short-term future channel prediction at latencies compatible with real-time V2X operation.
-Predictions that arrive too late are discarded to ensure consistency with the evolving physical system.
+Any predictions that arrive too late are discarded to maintain consistency with the evolving physical system.
 This integration exposes the key challenges of real-time NDT operation, including the computational cost of ray-traced propagation, the synchronization of mobility and geometry between platforms, and the scheduling required to deliver actionable predictions before the corresponding physical time occurs.
@@
   • We design and implement an end-to-end integration of VaN3Twin and the Tokyo Mobility DT into a latency-aware and predictive NDT;
-  • We introduce a timing-constrained workflow for future channel state prediction, where mobility updates, 3D scene synchronization, ray tracing, and network simulation are orchestrated to meet a real-time budget;
+  • We introduce a timing-constrained workflow for future channel state prediction, where mobility updates, 3D scene synchronization, ray tracing, and network simulation are orchestrated to meet a strict latency budget;
   • We demonstrate the integrated system in an urban testbed, quantifying the trade-offs between trajectory prediction error, channel prediction accuracy, and system latency.
@@
 Communication between system elements is structured according to their functional roles.
-A specific physical entity provides the interface between Physical Entities and the Mobility DT, leveraging a 2.4 GHz IEEE 802.11n Wi-Fi Control Plane (CP) for low latency over-the-air control operations.
+A specific physical entity provides the interface between physical entities and the Mobility DT, leveraging a 2.4 GHz IEEE 802.11n Wi-Fi Control Plane (CP) for low-latency over-the-air control operations.
 Meanwhile, the Mobility DT and VaN3Twin communicate over a wired Ethernet Local Area Network (LAN), reflecting their co-location at the network edge to minimize latency.
 All inter-component exchanges employ lightweight UDP messages with JSON-formatted payloads, enabling efficient and low-overhead transmission of state and control information.
-Message delivery is best-effort: occasional packet losses are tolerated and naturally overridden by subsequent updates, avoiding retransmission delays.
+Message delivery is best-effort: occasional packet losses are tolerated and naturally compensated for by subsequent updates, avoiding retransmission delays.
 All components rely on a shared time base, with timestamps carried in each message and clocks synchronized using standard network time synchronization mechanisms, ensuring consistent temporal alignment across mobility and communication predictions.
@@
 Based on the aggregated and synchronized mobility state, the Mobility DT computes short-term trajectory forecasts for all relevant dynamic objects.
-Network prediction requests toward VaN3Twin are then generated according to a configurable policy, which may be periodic or event-driven, and always target a single future time instant $t+h$.
+Network prediction requests to VaN3Twin are then generated according to a configurable policy, which may be periodic or event-driven, and always target a single future time instant $t+h$.
 The prediction horizon $h$ is chosen so as to remain compatible with the end-to-end latency budget; multi-horizon requests are also possible in principle but are outside the scope of this work.
-Upon receiving a request, VaN3Twin instantiates a future communication scenario corresponding to the target time $t+h$.
+Upon receiving a request, VaN3Twin generates a future communication scenario corresponding to the target time $t+h$.
 Predicted vehicle poses are used to construct a consistent 3D scene, which serves as input to the deterministic ray-based channel model formalized in Section III.
 The resulting channel realizations drive full-stack V2X simulations, possibly yielding protocol- and packet-level performance indicators.
-In parallel, the recently observed communication measurements are used to compensate for systematic prediction biases, such as offsets on the computed received power.
+In parallel, the recently observed communication measurements are used to compensate for systematic prediction biases, such as offsets in the computed received power.
 The predicted communication state is then returned to the Mobility DT and made available for network-level decision making.
@@
 This section provides a formal characterization of the integration presented in Section II.
-We describe the trajectory prediction model, formalize the ray-based channel prediction process, and introduce a compact notion of model complexity that links physical-model fidelity to end-to-end latency and feasible prediction horizons.
+We describe the trajectory prediction model, formalize the ray-based channel prediction process, and introduce a concise characterization of model complexity that links physical-model fidelity to end-to-end latency and feasible prediction horizons.
 III-A Mobility and Channel Prediction
@@
 The state evolution assumes a constant-velocity motion model, with $\mathbf{F}^{i}_{t}$ defined accordingly based on the sampling interval.
-$\mathbf{z}^{i}_{t}$ represents the state measurement vector, which the DT fuses into a unified state observation.
+$\mathbf{z}^{i}_{t}$ represents the state measurement vector, which the DT fuses into a unified state estimate.
 The temporal evolution of each vehicle follows a linear state-space model defined by the state-transition matrix $\mathbf{F}^{i}_{t}$, the control-input matrix $\mathbf{B}^{i}_{t}$, and the measurement model matrix $\mathbf{M}^{i}_{t}$, with process noise covariance $\mathbf{Q}^{i}_{t}$ and measurement noise covariance $\mathbf{R}^{i}_{t}$.
@@
 Each ray propagates through the environment and may undergo up to $N_{\mathrm{int}}$ interactions with surrounding surfaces.
-A ray reaching the antenna of vehicle $j$ constitutes a valid propagation path and is added to the set $\mathcal{P}^{i,j}_{\mathrm{valid}}$.
+A ray reaching the antenna of vehicle $j$ is part of a valid propagation path that is added to the set $\mathcal{P}^{i,j}_{\mathrm{valid}}$.
 Each path $p\in\mathcal{P}^{i,j}_{\mathrm{valid}}$ is characterized by its length $d_{p}^{i,j}$, delay $\tau_{p}^{i,j}=d_{p}^{i,j}/c$, angles of departure $(\theta^{(p)}_{\mathrm{T}},\varphi^{(p)}_{\mathrm{T}})$ and arrival $(\theta^{(p)}_{\mathrm{R}},\varphi^{(p)}_{\mathrm{R}})$, and by an ordered sequence of $N_{p}$ interactions.
@@
 Diffraction $\times$ $\times$ $\times$ $\times$ $\checkmark$
-Let $\tau_{e2e}$ be the total time elapsed between the acquisition of a measurement by a Physical Entity and the completion of the corresponding DT update. Formally:
+Let $\tau_{e2e}$ be the total time elapsed between the acquisition of a measurement by a vehicle and the completion of the corresponding DT update. Formally:
 $\tau_{e2e}=\tau_{m}+\tau_{w}+\tau_{tp}+\tau_{req}+\tau_{rt}+\tau_{req}+\tau_{w},$ (4)
 where $\tau_{m}$ is the measurement acquisition time, $\tau_{w}$ the communication delay between vehicles and the Mobility DT, $\tau_{tp}$ the Mobility DT update and trajectory prediction time, $\tau_{req}$ the inter-process communication delay, and $\tau_{rt}$ the execution time of the physical-modeling stage.
-Given the computational complexity of the ray tracing operation, $\tau_{rt}$ is the dominant and most variable value.
+Given the computational complexity of the ray tracing operation, $\tau_{rt}$ is the dominant and most variable component.
 Specifically, it depends on the level of physical detail included in the propagation model.
 To describe these configurations in a compact and general manner, we introduce the Ray Tracing Detail Index, denoted DI, a scalar metric used to classify the fidelity level of the physical-modeling stage.
-Each DI level corresponds to a specific configuration of the physical model, ranging from lightweight approximations to high-fidelity reconstructions.
+Each DI level corresponds to a specific configuration of the physical model, ranging from lightweight approximations to high-fidelity representations.
 Table I reports the five DI levels considered in this work.
-Lower levels include only the essential components needed for coarse propagation estimates, whereas higher levels progressively incorporate more detailed interaction mechanisms, finer spatial resolution, and deeper multipath exploration.
-Although in our implementation these fidelity levels map to different configurations of a ray-based propagation model, the abstraction provided by DI remains general and applicable to any physical-modeling backend supporting multiple complexity tiers.
+Lower levels include only the essential components needed for coarse propagation estimates, whereas higher levels progressively incorporate more detailed interaction mechanisms, finer spatial resolution, and higher-order multipath modeling.
+Although in our implementation these fidelity levels map to different configurations of a ray-based propagation model, the abstraction provided by DI remains general and applicable to any physical-modeling backend that supports multiple complexity tiers.
 As the DI increases, the execution time of the physical model, denoted $\tau_{rt}$, grows accordingly.
-To ensure causal consistency, the prediction horizon $h$ must exceed the end-to-end latency $\tau_{e2e}$.
+To ensure causality, the prediction horizon $h$ must exceed the end-to-end latency $\tau_{e2e}$.
 Increasing model fidelity therefore requires larger horizons, which in turn amplify trajectory prediction errors and their impact on channel estimation, revealing a fundamental trade-off between physical-model complexity, latency, and predictive accuracy.
@@
 IV Experimental and Simulated Results
-To demonstrate the practical operation of the proposed system, we consider a proof-of-concept deployment carried out in the Ookayama campus of the Institute of Science Tokyo, Japan complete with measured end-to-end latencies.
+To demonstrate the practical operation of the proposed system, we consider a proof-of-concept deployment carried out in the Ookayama campus of the Institute of Science Tokyo, Japan including measured end-to-end latencies.
 To complement the experimental analysis, we then consider a simulation study performed in the identical geometric scenario in which we systematically vary the DI and introduce synthetic perturbations into the trajectory prediction module.
@@
 Carrier Frequency 60 GHz
-Buildings and Roads Material Concrete [iturp2040]
-Trees Material Wood [iturp2040]
-Cars Material Metal [iturp2040]
+Buildings and Roads Material Concrete [8]
+Trees Material Wood [8]
+Cars Material Metal [8]
 Type of Antennas Panasonic WiGig RSU
@@
 Digital Model DI DI 5 (see Tab. I)
-The evaluation scenario involves three connected vehicles (Figure 3(a)) that interact with the Mobility DT through a $2.4$ GHz Wi-Fi CP, and are each equipped with a roof-mounted Panasonic WiGig RSU forming the $60$ GHz User Plane (UP).
-The Mobility DT and the VaN3Twin engine are deployed at the edge—corresponding to the location of Vehicle 3—and exchange data through a wired Ethernet LAN.
+The evaluation scenario involves three connected vehicles (Figure 3(a)) that interact with the Mobility DT through a $2.4$ GHz Wi-Fi CP, and are each equipped with a roof-mounted Panasonic WiGig RSU forming the $60$ GHz high-capacity User Plane (UP).
+The Mobility DT and VaN3Twin are deployed at the edge—corresponding to the location of Vehicle 3—and exchange data through a wired Ethernet LAN.
 Figure 3(b) illustrates the corresponding VaN3Twin ray tracing scenario of the considered street in Ookayama, Tokyo, Japan.
@@
 During the experimental campaign, the vehicles perform iperf-based data transmissions over the WiGig UP.
-Vehicle 1 (a Toyota COMS) and Vehicle 3 (a motorized wheel cart) remain stationary and maintain a direct Line of Sight (LoS) link, while Vehicle 2 (a Toyota Estima) follows a predefined trajectory (Figure 3(b)) that intermittently obstructs the link.
+Vehicle 1 (a Toyota COMS) and Vehicle 3 (a motorized wheel cart) remain stationary and maintain a direct Line of Sight (LoS) link, while Vehicle 2 (a Toyota Estima) follows a predefined trajectory (Figure 3(b)) that intermittently blocks the LoS path.
 The Mobility DT is used to predict, in real time, the $60$ GHz channel between Vehicles 1 and 3 via VaN3Twin, based on the predicted trajectory of Vehicle 2 with prediction horizon $h=500$ ms.
+Such predictions may be employed, for example, for proactive network topology management.
 The Mobility DT simultaneously collects from each vehicle its current position and the field-measured Received Signal Strength Indicators (RSSIs), which are employed for in-the-loop calibration of the prediction process.
-After computation, the predicted RSSI values are transmitted to the vehicles through the Wi-Fi CP.
+After computation, a control message is sent to the vehicles through the Wi-Fi CP.
 To quantify the effectiveness of this real-time predictive loop, we evaluate the accuracy of the RSSI forecasts produced by VaN3Twin along the trajectory of Vehicle 2.
@@
 Results show a maximum average RSSI prediction error of $1.01$ dB and a 95th-percentile error of $3.05$ dB, demonstrating accurate prediction of the $60$ GHz link dynamics under mobility-induced obstruction.
-Since VaN3Twin operates on predicted positions, channel accuracy is jointly limited by trajectory error and by the ray tracing DI, which directly determines the end-to-end latency and thus the minimum prediction horizon required for real-time operation.
+Since VaN3Twin operates on predicted positions, channel accuracy is jointly limited by trajectory prediction error and by the selected ray tracing DI, which directly determines the end-to-end latency and thus the minimum prediction horizon required for real-time operation.
 IV-B End-to-End System Latency
 Building on the formal definition introduced in Section III-B, this section reports on-field measurements of the end-to-end latency $\tau_{\mathrm{e2e}}$, decomposed according to (4).
-All the measured fixed latency components are summarized in Table III, while the resulting end-to-end latency distributions as a function of the DI level are shown in Figure 4.
-Fixed system latency components: measurement acquisition ($\tau_{m}$) was evaluated by collecting location and channel samples over $60$ s with a $100$ ms sampling interval, yielding approximately $600$ samples per experiment.
+Fixed system latency components: are summarized in Table III and constitute a fixed system latency budget, independent of the DI configuration.
+Measurement acquisition ($\tau_{m}$) was evaluated by collecting location and channel samples over $60$ s with a $100$ ms sampling interval, yielding approximately $600$ samples per experiment.
 No statistically significant differences were observed between static and dynamic conditions.
@@
 Mobility DT update time and trajectory prediction latency ($\tau_{tp}$) was obtained by profiling the time from the scenario update up to the execution of the linear prediction algorithm described in Section III-A.
-These components—resumed in Table III—constitute a fixed system latency budget, independent of the DI configuration.
 Ray tracing execution times ($\tau_{rt}$): were measured for the five DI configurations defined in Table I using VaN3Twin with Sionna RT v1.2.1.
@@
 End-to-end latency $\tau_{\mathrm{e2e}}$: is obtained by combining the fixed system latency components with the DI-dependent ray tracing execution time.
-The resulting distributions of $\tau_{\mathrm{e2e}}$ for each DI configuration are reported in Figure 4, highlighting the progressive increase in both latency and variability as the ray tracing complexity grows.
+The distributions of $\tau_{\mathrm{e2e}}$ in Figure 4 show increasing latency and variability with ray tracing complexity, peaking at an average $250$ ms for DI 5.
 These results demonstrate that improving channel modeling fidelity by increasing the DI level inevitably enlarges the end-to-end latency budget.
@@
 IV-C Impact of trajectory prediction error
-Because the end-to-end latency of the DT determines how far into the future vehicle positions must be predicted, longer latencies inevitably amplify trajectory prediction errors.
+Because the end-to-end latency of the DT determines how far into the future vehicle positions must be predicted, longer latencies inevitably exacerbate trajectory prediction errors.
 This section quantifies how such latency-induced positional inaccuracies propagate into ray-traced channel prediction.
-Let $\mathbf{p}(t)=[x(t),\,y(t)]^{\top}$ denote the ground-truth position at time $t$, and let $\tilde{\mathbf{p}}(t)=\mathbf{p}(t)+\mathbf{e}(t)$, be its perturbed counterpart, where $\mathbf{e}(t)$ models the prediction error.
+Let $\mathbf{p}(t)=[x(t),\,y(t)]^{\top}$ denote the ground-truth position at time $t$, and $\tilde{\mathbf{p}}(t)=\mathbf{p}(t)+\mathbf{e}(t)$ its perturbed counterpart, where $\mathbf{e}(t)$ models the prediction error.
 A normalized parameter $k\in[0,1]$ controls the maximum error magnitude through $\varepsilon_{k}=k\,\varepsilon_{\max}$, with $\varepsilon_{\max}=1$ m in the considered setup.
@@
 Figure 5 shows the dependence of the computed RMSE (blue) on the perturbation level $k$, focusing on measurements collected during the LoS to non-LoS transition caused by the obstructing vehicle (Vehicle 2).
-The relationship is clearly non-linear, with a rapid error increase for small positional deviations.
-For the considered scenario, $k\approx 0.1$ already yields an RMSE of about $20$ dB on RSSI prediction.
-Since small geometric inaccuracies already induce large prediction errors, these results also anticipate the sensitivity of the LoS classification, analyzed next.
+The relationship is clearly non-linear, with a rapid increase in error for small positional deviations.
+For the considered scenario, $k\approx 0.1$ already yields an RMSE of approximately $20$ dB in the RSSI prediction.
+Since small geometric inaccuracies already induce large prediction errors, these results also foreshadow the sensitivity of the LoS classification, analyzed next.
 LoS Classification Robustness: trajectory perturbations may also change the geometric visibility between a transmitter and a receiver.
@@
 V Conclusions
-In this work, we presented and experimentally validated a real-time, predictive NDT for V2X communications, explicitly designed around end-to-end latency constraints.
+In this work we presented and experimentally validated a real-time, predictive NDT for V2X communications, explicitly designed around end-to-end latency constraints.
 By integrating a live Mobility Digital Twin with the VaN3Twin engine, we implemented a closed-loop workflow capable of operating in synchrony with the physical environment.
@@
 Overall, this work shows that real-time integration of mobility Digital Twins and ray-tracing-based network modeling is feasible and practically deployable.
-Future work will extend this latency-aware framework toward real-time prediction of higher-layer network performance metrics leveraging on VaN3Twin full-stack V2X simulation capability.
+Future work will extend this latency-aware framework toward real-time prediction of higher-layer network performance metrics leveraging VaN3Twin full-stack V2X simulation capability.
 Acknowledgment
 This work was partially supported by the European Union - Next Generation EU under the Italian National Recovery and Resilience Plan (NRRP), Mission 4, Component 2, Investment 1.3, CUP D43C22003080001, partnership on “Telecommunications of the Future” (PE00000001 - program “RESTART”).
+This work was also supported by Japan Science and Technology Agency (JST) as part of Adopting Sustainable Partnerships for Innovative Research Ecosystem (ASPIRE), Grant Number JPMJAP2517.